# SASSL: Enhancing Self-Supervised Learning via Neural Style Transfer

### v2

@@ Abstract
Self-supervised learning relies heavily on data augmentation to extract meaningful representations from unlabeled images. While existing state-of-the-art augmentation pipelines incorporate a wide range of primitive transformations, these often disregard natural image structure. Thus, augmented samples can exhibit degraded semantic information and low stylistic diversity, affecting downstream performance of self-supervised representations. To overcome this, we propose SASSL: Style Augmentations for Self Supervised Learning, a novel augmentation technique based on Neural Style Transfer. The method decouples semantic and stylistic attributes in images and applies transformations exclusively to the style while preserving content, generating diverse augmented samples that better retain their semantic properties. Experimental results show our technique achieves a top-1 classification performance improvement of more than $2\%$ on ImageNet compared to the well-established MoCo v2. We also measure transfer learning performance across five diverse datasets, observing significant improvements of up to $3.75\%$. Our experiments indicate that decoupling style from content information and transferring style across datasets to diversify augmentations can significantly improve downstream performance of self-supervised representations.
@@ 1 Introduction
Data labelling is a challenging and expensive process, which often serves as a barrier to build machine learning models to solve real-world problems. Self-supervised learning (SSL) is an emerging machine learning paradigm that helps to alleviate the challenges of data labelling, by using large corpora of unlabeled data to pre-train models to learn robust and general representations. These representations can be efficiently transferred to downstream tasks, resulting in performant models which can be constructed without access to large pools of labeled data. SSL methods have shown promising results in recent years, matching and in some cases exceeding the performance of bespoke supervised models with small amounts of labelled data.
Given the lack of labels, SSL relies on pretext tasks, i.e., predefined tasks where pseudo-labels can be generated. Some examples include contrastive learning (Chen et al., 2020a; He et al., 2020), clustering (Caron et al., 2021; 2020; Assran et al., 2022), and generative modeling (He et al., 2022; Devlin et al., 2018). Many of these pretext tasks involve training the model to distinguish between different views of the same input and inputs corresponding to different samples. For these tasks, the way input data is augmented is crucial for the network to learn useful invariances and extract robust representations (Chen et al., 2020a). While state-of-the-art augmentations incorporate a wide range of primitive color, spectral and spatial transformations, they often disregard the natural structure of an image. As a result, SSL pre-training methods may generate augmented samples with degraded semantic information, and may be less able to capture diverse visual attributes.
To overcome this limitation, we propose Style Augmentations for Self Supervised Learning (SASSL), a novel SSL data augmentation technique based on Neural Style Transfer to obtain stronger augmented samples via semantically-aware preprocessing. In contrast to handcrafted augmentation techniques operating on specific formats (e.g. pixel or spectral domain), SASSL disentangles an image into perceptual (style) and semantic (content) representations that are learned from data. By applying transformations exclusively to the style of an image while preserving its content, we can generate augmented samples with diverse styles that retain the original semantic properties.
@@
  • We propose SASSL, a novel data augmentation technique based on Neural Style Transfer that naturally preserves semantic information while varying stylistic information (Section 4).
  • We empirically demonstrate improved downstream task performance by incorporating our method into MoCo v2, without additional hyperparameter tuning. We report more than $2\%$ top-1 accuracy improvement on the challenging ImageNet dataset (Section 5.1).
  • We show our augmentation method learns more robust representations by measuring their transfer learning capabilities on five diverse datasets. SASSL improves linear probing performance by up to $3.75\%$ and fine-tuning by up to $1\%$ on out-of-distribution tasks (Section 5.2).
  • We observe that balancing stylization using representation blending and image interpolation performs best, and adding external style datasets for transfer can improve performance (Section 5.3).
(a) Proposed data augmentation pipeline.
@@
Figure 1: Towards diverse SSL data augmentation via Neural Style Transfer. We propose SASSL, a novel augmentation method based on Style Transfer that generates semantic-aware transformations of pre-training images by operating exclusively on their style. Given a pre-training image and an external style representation, SASSL combines the content attributes of the image with the external style, producing an augmented view that better preserves semantic attributes. By controlling the stylization strength via feature and pixel interpolations, we empirically show SASSL promotes robust representations by incorporating Style Transfer to default SSL augmentation pipelines.
@@ 2.1 Data Augmentation in SSL
Typical data augmentation methods in SSL applied to vision tasks correspond to image cropping and resizing, flipping, rotation, color augmentation, noise addition, and solarization. Examples of methods using these are MoCo (He et al., 2020), SimCLR (Chen et al., 2020a), BYOL (Grill et al., 2020), and SimSiam (Chen & He, 2021), among others. Other research has explored how to select augmentation hyperparameters to learn more robust and general features (Wagner et al., 2022; Reed et al., 2021; Tian et al., 2020). In contrast, our method proposes a new preprocessing technique that can be incorporated in any existing augmentation pipeline, obtaining improved performance without additional hyperparameter tuning.
Novel augmentation techniques have been proposed for both supervised and self-supervised learning. For example, Han et al. (2022) propose a novel method to apply multiple data augmentation techniques in a local fashion by splitting the image into patches. This is aligned with other local augmentation techniques such as random erasing (Zhong et al., 2020), CutMix (Yun et al., 2019), Mixup (Zhang et al., 2017), among others. While these methods focus on masking or severely distorting image patches, disregarding their semantic information, our approach is a global transformation that decouples an image into its appearance and semantic attributes. This allows applying transformations exclusively to its appearance without affecting the image semantics.
@@ 2.2 Neural Style Transfer
Recent image generation and style transfer algorithms (Liu et al., 2021; Heitz et al., 2021; Jing et al., 2020; Yoo et al., 2019; Risser et al., 2017; Gatys et al., 2017; Chen et al., 2021a; Wang et al., 2020) use deep learning, such as CNNs, to measure the texture similarity between two images in terms of the distance between their neural activations. Specifically, feature maps of a pre-trained classifier such as VGG-19 (Simonyan & Zisserman, 2014) are extracted and low-order moments of their distribution (Kessy et al., 2018; Huang & Belongie, 2017; Sheng et al., 2018) are used as texture descriptors. By matching such feature statistics, CNN-based techniques have shown promising results transferring texture between arbitrary images, significantly improving over classic texture synthesis approaches (Portilla & Simoncelli, 2000; Zhu et al., 2000; Heeger & Bergen, 1995).
A large body of work in CNN-based style transfer focuses on artistic applications such as reproducing an artwork style and imposing it over a scene of interest. These methods adopt either (i) an iterative optimization approach (Heitz et al., 2021; Risser et al., 2017; Gatys et al., 2016; 2017; Li et al., 2017a), where an initial guess is gradually transformed to depict a style of interest or (ii) an encoding-decoding strategy (Liu et al., 2021; Jing et al., 2020; Li et al., 2017b; Chen et al., 2021a; Wang et al., 2020), where one or more CNN-based image generators are trained to impose a target texture over the scene of interest in a single forward pass. While selecting one of these implies a trade-off between synthesis quality and computational cost, in most cases the generated stylization shows an unnatural appearance, i.e., it does not capture most of the texture details of a style image.
Alternative techniques show impressive style generalization with much simpler transformations. In their seminal work, Dumoulin et al. (2017) propose a style transfer technique that extracts a low-dimensional representation from an arbitrary style image. This representation is then used to predict scaling and shifting values that are used to normalize the feature maps of a content image. The authors empirically show that, by learning such a contracted representation of style images, their method can represent styles from a wide range of unseen domains, leading to a faithful stylization.
Other works have applied style transfer in context of supervised learning, to produce additional labelled data (Zheng et al., 2019), remove model biases (Geirhos et al., 2018) and improve robustness against adversarial attacks (Hong et al., 2021). To our knowledge, this is the first work using neural style transfer to learn more performant self-supervised representations, without labelled data.
@@ 3.1 Self-Supervised Learning
Our proposed work deals with data augmentation in SSL. More precisely, we propose a novel preprocessing operation to enhance representation learning. Before introducing our work, we review the fundamentals of SSL by revisiting a well established technique: SimCLR.
A Simple Framework for Contrastive Learning. SimCLR learns compressed representations by maximizing the agreement between differently augmented views of the same data example in a latent space, while minimizing the similarity to different examples. Potential augmentations include random cropping, flipping, color jitter, blurring, and solarization. By maximizing the similarity of augmented training samples, the network learns to create robust representations that are invariant to simple distortions that could occur in the real world, and which should not affect the semantic content of an image. By simultaneously contrasting against other examples, the representations learn to extract information that is unique to each example. The combination of the two leads to the extraction of useful semantic information from the data.
Given a batch of $N$ input images $\{{\bm{i}}_{k}\}_{k=1}^{N}$, $2N$ augmented samples are generated by applying distinct transformations to each image. These transformations correspond to the same data augmentation pipeline. Let $R$ correspond to all possible augmentations. Then, positive pairs correspond to augmented views of the same input sample, and negative pairs correspond to views coming from different input images. Based on this, the $2N$ augmented samples $\{\tilde{{\bm{i}}}_{l}\}_{l=1}^{2N}$ can be organized so that indices $l=2k-1$ and $l=2k$ correspond to views of the same input sample
@@
Once augmented views are obtained, a representation is computed using an encoder (typically a CNN-based model). The representations are then fed to a projection head which further compresses the representation into a lower-dimensional manifold where different views of the same image are close together and those from different images are far apart. Let $h$ and $g$ be the encoder (e.g. a ResNet-50 backbone) and projection head (e.g. a multilayer perceptron), respectively. Then, embeddings are obtained for each augmented sample
${\bm{z}}_{l}=g\circ h(\tilde{{\bm{i}}}_{l})$ (2)
SimCLR relies on the normalized temperature-scaled cross entropy loss (NT-Xent) to learn how to distinguish between positive pairs of augmented samples. First, the cosine similarity distance between every pair of embeddings is computed
@@
The final step is to use a contrastive loss to train the model. The contrastive loss compares the embeddings of positives and encourages them to be as similar as possible in the latent space. Since the loss is normalized, it naturally forces the representations of views from two different images (negatives) to be far away from each other.
$\mathcal{L}=$ $\ \frac{1}{2N}\sum_{k=1}^{N}\big{[}l(2k-1,2k)+l(2k,2k-1)\big{]}$ (4)
$l(m,n)=$ $\ -\log\bigg{(}\frac{\exp(s_{m,n}/\tau)}{\sum_{l=1}^{2N}\mathbbm{1}_{m\neq n}\exp(s_{m,l}/\tau)}\bigg{)}$ (5)
where $\tau\in\mathbb{R}_{++}$ is the temperature factor and $\mathbbm{1}$ the indicator function. While SimCLR is a simple framework, it pushed the state-of-the-art significantly on a wide range of downstream tasks including image classification, object detection, and semantic segmentation. Follow up works to SimCLR such as Grill et al. (2020); Chen & He (2021); Caron et al. (2020; 2021); Assran et al. (2022); Zbontar et al. (2021); Chen et al. (2020b; 2021b); Bardes et al. (2021) have proposed modifications to this setup which attempt to further improve the downstream task performance.
@@ 3.2 Neural Style Transfer
Neural Style Transfer algorithms aim to preserve the content of an image while imposing the style of another. These approaches assume that the statistics of shallower layers of a trained CNN encode style, while deeper layers encode content. Thus, early Style transfer algorithms worked by passing the pair of content and style images to a trained image encoder and optimizing the content image to produce activations which had similar statistics to the style image at shallower layers while keeping the deeper layer activations unchanged (Simonyan & Zisserman, 2014).
Once semantic and texture features are extracted, a stylized image with the semantic attributes of the content image and the texture properties of the style image is generated. The two main approaches are image-optimization based and model-based methods. In what follows, we introduce a model-based Style Transfer technique adopted by our proposed method due on its generalization and efficiency properties. Jing et al. (2019) provide an in-depth study of image stylization techniques.
Fast Style Transfer. Dumoulin et al. (2017) proposed an arbitrary style transfer method with remarkable generalization properties. Their algorithm, Fast Style Transfer, has been shown to accurately represent unseen artistic styles by training a model to predict first and second moments of latent image representations at multiple scales. Such moments are used as arguments of a special form of instance normalization, denominated conditional instance normalization (CIN), to impose style over arbitrary input images. Below we formally introduce the Fast Style Transfer algorithm.
Given a content image ${\bm{i}}_{c}\in\mathbb{R}^{C\times H_{c}\times W_{c}}$ and a style image ${\bm{i}}_{s}\in\mathbb{R}^{C\times H_{s}\times W_{s}}$, the produced stylized image ${\bm{i}}_{\text{cs}}$ corresponds to
@@
where $\mathcal{T}$ is a stylization network and ${\bm{z}}_{s}$ is a feature embedding extracted from the style image using a pre-trained feature extractor denoted as $\mathcal{F}$, e.g., InceptionV3 (Szegedy et al., 2016).
${\bm{z}}_{s}=$ $\ \mathcal{F}({\bm{i}}_{s})\in\mathbb{R}^{D}$ (7)
Here, we assume ${\bm{z}}_{s}$ to be a contracted embedding of the style image ($D\ll CH_{s}W_{s}$). The stylization network $\mathcal{T}$ is comprised by $L$ blocks $\{t_{l}\}_{l=1}^{L}$. These extract high-level features from the content image, align them to the style embedding ${\bm{z}}_{s}$ and map the resulting features to the pixel domain.
The style of ${\bm{i}}_{s}$ encapsulated in ${\bm{z}}_{s}$ is transferred to the content image using CIN. This is applied to a particular set of layers to impose the target texture and color scheme by aligning feature maps at different scales. We define the set of layers where CIN is applied as $\mathcal{A}$. Conditional instance normalization (CIN) consists of an extended form of instance normalization where the target mean and standard deviation are extracted from an arbitrary style representation ${\bm{z}}$. Given an input ${\bm{i}}\in\mathbb{R}^{C\times H\times W}$ and a style representation ${\bm{z}}\in\mathbb{R}^{D}$, CIN is defined as
$\hat{{\bm{i}}}=$ $\ \text{CIN}({\bm{i}},{\bm{z}})$ (8)
$\hat{{\bm{i}}}^{(k)}=$ $\ \gamma^{(k)}({\bm{z}})\bigg{(}\frac{{\bm{i}}^{(k)}-\mathbb{E}[{\bm{i}}^{(k)}]}{\sigma({\bm{i}}^{(k)})}\bigg{)}+\lambda^{(k)}({\bm{z}}),\ k\in\{1,\dots,C\}$ (9)
where ${\bm{i}}^{(k)}$ corresponds to the $k$-th input channel, and the sample mean and standard deviation are computed along its spatial support. Here, $\gamma^{(k)},\lambda^{(k)}:\mathbb{R}^{D}\mapsto\mathbb{R}$ are trainable functions that predict scaling and offset values from the latent representation ${\bm{z}}$ for the $k$-th input channel. Following this, the layers in the stylization network are characterized by
@@
where the network input corresponds to $\bm{\phi}_{c,0}={\bm{i}}_{c}$. We use the subscript $l$ in the CIN operation to indicate that each layer has its own $\gamma$ and $\lambda$ functions to normalize feature maps independently.
Our method uses the Fast Style Transfer algorithm. Given its generalization properties and low-dimensional style representations, it is a good match for our framework, where style representations from multiple domains must be extracted, manipulated and transferred in an efficient manner.

### v4

@@ Abstract
Existing data augmentation in self-supervised learning, while diverse, fails to preserve the inherent structure of natural images. This results in distorted augmented samples with compromised semantic information, ultimately impacting downstream performance. To overcome this limitation, we propose SASSL: Style Augmentations for Self Supervised Learning, a novel data augmentation technique based on Neural Style Transfer. SASSL decouples semantic and stylistic attributes in images and applies transformations exclusively to their style while preserving content, generating diverse samples that better retain semantic information. SASSL boosts top-1 image classification accuracy on ImageNet by up to 2 percentage points compared to established self-supervised methods like MoCo, SimCLR, and BYOL, while achieving superior transfer learning performance across various datasets. Because SASSL can be performed asynchronously as part of the data augmentation pipeline, these performance impacts can be obtained with no change in pretraining throughput.
@@ 1 Introduction
Data labelling is a challenging and expensive process, which often serves as a barrier to build machine learning models to solve real-world problems. Self-supervised learning (SSL) is an emerging machine learning paradigm that helps to alleviate the challenges of data labelling, by using large corpora of unlabeled data to pretrain models to learn robust and general representations. These representations can be efficiently transferred to downstream tasks, resulting in performant models which can be constructed without access to large pools of labeled data. SSL methods have shown promising results in recent years, matching and in some cases exceeding the performance of bespoke supervised models with small amounts of labelled data.
Given the lack of labels, SSL relies on pretext tasks, i.e., predefined tasks where pseudo-labels can be generated. These include contrastive learning (Chen et al., 2020a; He et al., 2020), clustering (Caron et al., 2021; 2020; Assran et al., 2022), and generative modeling (He et al., 2022; Devlin et al., 2018). Many pretext tasks involve training the model to distinguish between different views of the same input and inputs corresponding to different samples. For these tasks, the way input data is augmented is crucial to learn useful invariances and extract robust representations (Chen et al., 2020a). While state-of-the-art augmentations incorporate a wide range of color, spectral and spatial transformations, they often disregard the natural structure of an image. As a result, SSL pretraining methods may generate augmented samples with degraded semantic information, and may be less able to capture diverse visual attributes.
To tackle this challenge, we propose Style Augmentations for Self Supervised Learning (SASSL), a novel SSL data augmentation technique based on Neural Style Transfer to generate semantically consistent augmented samples. In contrast to augmentation techniques operating on specific formats (e.g. pixel or spectral domain), SASSL disentangles an image into perceptual (style) and semantic (content) representations that are learned from data. Applying transformations only to the style of an image while preserving its content, we can generate images with diverse appearance that retain the original semantic properties.
@@
  • We propose SASSL, a novel data augmentation technique based on Style Transfer that naturally preserves semantic properties while diversifying style (Section 4).
  • We empirically show improved downstream performance on ImageNet Deng et al. (2009) by incorporating SASSL in methods such as MoCo, BYOL and SimCLR without hyperparameter tuning (Sections 5.1, 5.4).
  • We show SASSL learns stronger representations by measuring their transfer learning capabilities on various datasets. Our method boosts linear probing performance by up to $10\%$ and fine-tuning by up to $6\%$ on out-of-distribution tasks (Section 5.2).
(a) Data augmentation pipeline.
@@
Figure 1: Towards diverse SSL data augmentation via Neural Style Transfer. We propose SASSL, a novel augmentation technique that leverages Style Transfer to create pretraining views that are semantically aware, focusing solely on modifying the image’s appearance while preserving its content. SASSL combines the image’s content with the texture of an external reference style, generating augmented views that better retain the image’s semantic meaning. By incorporating Style Transfer into traditional SSL augmentation pipelines and controlling the stylization strength through gradual blending of style features and pixel values, SASSL promotes stronger representations compared to well-established SSL methods.
@@ 2.1 Data Augmentation in SSL
Typical data augmentation methods applied to vision tasks include image cropping and resizing, flipping, rotation, color augmentation, noise addition, and solarization. Examples of methods using these are MoCo (He et al., 2020), SimCLR (Chen et al., 2020a), BYOL (Grill et al., 2020), and SimSiam (Chen & He, 2021), among others. Other work (Caron et al., 2020) shows how generating additional augmentations using this strategy can improve performance relative to two view approaches, though this strategy decreases throughput and requires additional compute. Other research explore how to select augmentation hyperparameters to learn more robust and general features (Wagner et al., 2022; Reed et al., 2021; Tian et al., 2020). In contrast, SASSL proposes a new preprocessing technique that can be incorporated in existing augmentation pipelines, boosting performance without additional hyperparameter tuning or changes in pretraining throughput.
Previous SSL work has explored semantic-aware augmentation approaches. Purushwalkam & Gupta (2020) leverage natural video transformations occuring in videos as an alternative to learning from object-centric datasets. Lee et al. (2021) introduces an auxiliary loss to capture the difference between augmented views, leading to better performance on tasks where semantic information is lost due to aggressive augmentation. Bai et al. (2022) propose an alternative augmentation pipeline to prevent loss of semantics by gradually increasing the strength of augmentations. While these methods modify the pretraining loss and require keeping track of augmentation hyperparameters, SASSL integrates seamlessly into existing pipelines without additional loss terms or auxiliary data. Our method complements default data augmentation pipelines with a content-preserving transformation to obtain stronger image representations.
@@ 2.2 Neural Style Transfer
Recent image generation and Style Transfer algorithms (Liu et al., 2021; Heitz et al., 2021; Jing et al., 2020; Yoo et al., 2019; Risser et al., 2017; Gatys et al., 2017; Chen et al., 2021a) use CNNs to measure the texture similarity between two images in terms of the distance between their neural activations. Feature maps of a pretrained classifier such as VGG-19 (Simonyan & Zisserman, 2014) are extracted and low-order moments of their distribution (Kessy et al., 2018; Huang & Belongie, 2017; Sheng et al., 2018) are used as texture descriptors. By matching such feature statistics, these techniques have shown promising results transferring texture between arbitrary images, improving over classic texture synthesis approaches (Portilla & Simoncelli, 2000; Zhu et al., 2000; Heeger & Bergen, 1995).
A large body of work focuses on artistic applications, reproducing an artwork style over a scene of interest. These methods adopt either (i) an iterative optimization approach (Risser et al., 2017; Gatys et al., 2016; 2017; Li et al., 2017a), where an initial guess is gradually transformed to depict a style of interest or (ii) an autoencoding approach (Liu et al., 2021; Li et al., 2017b; Chen et al., 2021a; Wang et al., 2020), where one or more CNN image generators are trained to impose a target texture in a single forward pass. While selecting an approach implies a trade-off between synthesis quality and computational cost, in both cases the generated stylization shows an unnatural appearance, i.e., it often lacks the qualities of a real-world scene.
In the context of data augmentation for supervised learning, Geirhos et al. (2018) and Zheng et al. (2019) addressed texture bias and generated training samples via pre-stylized datasets or stylizing using a small collection of style images. Hong et al. (2021) explored Style Transfer to improve robustness against adversarial attacks. Jackson et al. (2019) incorporated Style Transfer as a transformation in the augmentation pipeline. While their approach of randomly mixing content and style representations yield promising results, it neglects potential distortions introduced by the Style Transfer network bottleneck. SASSL, in contrast, integrates Style Transfer in a self-supervised setting. Our approach generates diverse augmentations via either pre-computed style representations from external datasets or in-batch stylization with training samples as style references. Importantly, SASSL preserves semantic information through pixel interpolation and feature blending, mitigating the loss of details inherent in Style Transfer networks.
@@ 3.1 Self-Supervised Learning
Traditional SSL methods learn compressed representations by maximizing the agreement between differently augmented views of the same data example in a latent space. They do this following the template originally proposed by SimCLR (Chen et al., 2020a). In this setup the input is split into multiple views using data augmentations, encoded into a representation, and then further projected into an embedding over which the loss is computed. There are many potential augmentations that can be used, including (but not limited to) random cropping, flipping, color jitter, blurring, and solarization. By maximizing the similarity of augmented training samples, the network learns to create robust representations that separate meaningful semantic content from simple distortions that could occur in the real world, and which should not affect the semantic content of an image.
Given a batch of $N$ input images $\{{\bm{i}}_{k}\}_{k=1}^{N}$, $2N$ augmented samples are generated by applying distinct transformations to each image. These transformations correspond to the same data augmentation pipeline. Let $R$ correspond to all possible augmentations. Then, positive pairs correspond to augmented views of the same input sample, and negative pairs correspond to views coming from different input images. Based on this, the $2N$ augmented samples $\{\tilde{{\bm{i}}}_{l}\}_{l=1}^{2N}$ can be organized so that indices $l=2k-1$ and $l=2k$ correspond to views of the $k$-th input sample
@@
Once augmented views are obtained, a representation is computed using an image encoder (typically a CNN model). The representations are then fed to a projection head which further compresses them into a lower-dimensional manifold where different views of the same image are close together and those from different images are far apart. Let $h$ and $g$ be the encoder (e.g. a ResNet-50 backbone) and projection head (e.g. an MLP layer), respectively. Then, embeddings are obtained for each augmented sample as ${\bm{z}}_{l}=g\circ h(\tilde{{\bm{i}}}_{l})$.
SimCLR uses the normalized temperature-scaled cross entropy loss (NT-Xent) to learn how to identify positive pairs of augmented samples. First, the cosine similarity of every pair of embeddings is computed
@@
The model is then trained using a contrastive loss by comparing the embeddings of positives, forcing them to be similar to each other. Since the loss is normalized, it naturally forces the representations of views from two different images (negatives) to be distant from each other.
$\mathcal{L}=$ $\ \frac{1}{2N}\sum_{k=1}^{N}\big{[}\ell(2k-1,2k)+\ell(2k,2k-1)\big{]},\quad\ell(m,n)=\ -\log\bigg{(}\frac{\exp(s_{m,n}/\tau)}{\sum_{l=1}^{2N}\mathbbm{1}_{m\neq n}\exp(s_{m,l}/\tau)}\bigg{)}$ (3)
where $\tau\in\mathbb{R}_{++}$ is the temperature factor and $\mathbbm{1}$ the indicator function. While SimCLR is a simple framework, it pushed the state-of-the-art significantly on a wide range of downstream tasks including image classification, object detection, and semantic segmentation.
Follow up works to SimCLR such as MoCo (Chen et al., 2020b; 2021b), BYOL (Grill et al., 2020) and SimSiam (Chen & He, 2021), among others (Caron et al., 2020; 2021; Assran et al., 2022; Zbontar et al., 2021; Bardes et al., 2021), have largely maintained this template, but have proposed modifications to this setup (e.g. new losses, architectures, or augmentation strategies) which attempt to further improve the downstream task performance.
@@ 3.2 Neural Style Transfer
Style Transfer techniques combine the semantics (content) of an image with the visual characteristics (style) of another image. These assume that the statistics of shallower layers of a trained CNN encode style, while deeper layers encode content. Seminal techniques are based on an optimization-based approach, passing a pair of content and style images to a CNN encoder and optimizing over a randomly initialized image to produce activations with similar statistics to the style image at shallower layers and similar activations to the content image at deeper ones (Gatys et al., 2015). This way, a stylized image is generated, comprising the semantic and texture attributes of interest.
While optimization-based methods generate a diverse stylization due to a random image initialization, autoencoding methods utilize an image decoder to efficiently stylize arbitrary image pairs on a single forward pass. In what follows, we introduce the autoencoding Style Transfer technique adopted by our proposed method due on its generalization and efficiency properties. For an in-depth survey of Style Transfer methods, refer to Jing et al. (2019).
Fast Style Transfer. Dumoulin et al. (2017) proposed an arbitrary Style Transfer method with remarkable generalization properties. Their algorithm, Fast Style Transfer, accurately represents unseen artistic styles by training a model to predict first and second moments of latent image representations at multiple scales. Such moments are used as arguments of a special form of instance normalization, denominated conditional instance normalization (CIN), to impose style over arbitrary input images.
Given a content image ${\bm{i}}_{c}\in\mathbb{R}^{C\times H_{c}\times W_{c}}$ and a style image ${\bm{i}}_{s}\in\mathbb{R}^{C\times H_{s}\times W_{s}}$, Fast Style Transfer produces a stylized image ${\bm{i}}_{\text{cs}}$ that corresponds to
@@
where $\mathcal{T}$ is a stylization network and ${\bm{z}}_{s}=\mathcal{F}({\bm{i}}_{s})\in\mathbb{R}^{D}$ is an embedding extracted from the style image via a feature extractor $\mathcal{F}$, e.g., InceptionV3 (Szegedy et al., 2016).
We assume ${\bm{z}}_{s}$ to be a contracted embedding of the style image ($D\ll CH_{s}W_{s}$). The stylization network $\mathcal{T}$ is comprised by $L$ blocks $\{t_{l}\}_{l=1}^{L}$. $\mathcal{T}$ extracts high-level features from the content image, aligns them to the style embedding ${\bm{z}}_{s}$ and maps the resulting features to the pixel domain. The style of ${\bm{i}}_{s}$ encapsulated in ${\bm{z}}_{s}$ is transferred to the content image using CIN. This is applied to a particular set of layers to impose the target texture and color scheme by aligning feature maps at different scales.
[…]
figureFeature blending and image interpolation. A fine-grained control over the final stylized image is obtained via interpolation and blending factors $\alpha$ and $\beta$ that operate in the feature and pixel domains. This prevents augmented views from losing semantic information due to strong transformations.
We define the set of layers where CIN is applied as $\mathcal{A}$. The normalization imposed via CIN consists of an extended form of instance normalization where the target mean and standard deviation are extracted from a style representation ${\bm{z}}$.
Given an input ${\bm{i}}\in\mathbb{R}^{C\times H\times W}$ and a style representation ${\bm{z}}\in\mathbb{R}^{D}$, CIN is defined as
$\hat{{\bm{i}}}=$ $\ \text{CIN}({\bm{i}},{\bm{z}})\in\mathbb{R}^{C\times H\times W}$ (5)
$\hat{{\bm{i}}}^{(k)}=$ $\ \gamma^{(k)}({\bm{z}})\bigg{(}\frac{{\bm{i}}^{(k)}-\mathbb{E}[{\bm{i}}^{(k)}]}{\sigma({\bm{i}}^{(k)})}\bigg{)}+\lambda^{(k)}({\bm{z}})$ (6)
where ${\bm{i}}^{(k)},\ k\in\{1,\dots,C\}$ corresponds to the $k$-th input channel, and the sample mean $\mathbb{E}[{\bm{i}}^{(k)}]$ and standard deviation $\sigma({\bm{i}}^{(k)})$ are computed along its spatial support. Here, $\gamma^{(k)},\lambda^{(k)}:\mathbb{R}^{D}\mapsto\mathbb{R}$ are trainable functions that predict scaling and offset values from the latent representation ${\bm{z}}$ for the $k$-th input channel. The layers in $\mathcal{T}$ are characterized by
@@
where the input of the stylization network corresponds to $\bm{\phi}_{c,0}={\bm{i}}_{c}$. The subscript $l$ in the CIN operation indicates that each layer has its own $\gamma$ and $\lambda$ functions to normalize features independently.
Our method uses the Fast Style Transfer algorithm. Given its generalization properties and low-dimensional style representations, it is a good match for our framework, where style representations from multiple domains must be efficiently extracted, manipulated and transferred.